OpenStudy (anonymous):

x-y=2. can someone show me how do this equation? thanks

OpenStudy (anonymous):

How do we know that? Let's see: (x+y)² = (x+y)(x+y) = x(x+y) + y(x+y) {distributive law) = x²+xy + xy+y² = x²+2xy+y²

OpenStudy (anonymous):

Deos that help

OpenStudy (anonymous):

no I still don't understand it. I need to graph it and th emath book we are using doesn't really show it.

OpenStudy (anonymous):

ok then ohld on plz

OpenStudy (anonymous):

There is a way to make this more obvious. To the left is a square. Each side is x+y. What is the area? Well, it is (x+y)^2. But, we can also find the area, by summing up the areas of the four rectangles (two of which are squares) which make up our big square: area = x^2+xy+xy+y^2 = x^2+2xy+y^2

OpenStudy (anonymous):

man it did not show the gragh

OpenStudy (anonymous):

i am beginning to hate this site. It keeps locking up or refreshing. Man is that a pain